# A note on Ramsey numbers for minors

### 11 Mar 2026 v1

@@ 3 Concluding remarks
We showed that $R_{h}(k;2)=ck\sqrt{\log k}(1+o(1))$, where $c$ is a constant, $1\leq c\leq 1.03$.... We believe that $c=1$. It would be interesting to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
To prove the lower bound on $R_{h}(k;\ell)$, one can also use the following packing result by taking $G$ being an $H$-factor, where $H$ is a graph on $k\sqrt{\log k}(1+o(1))$ vertices, density $\beta k\sqrt{\log k}$ and no $K_{k}$-minor. Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ own color. Decompose the remaining edges of $K_{n}$ into forests using Nash-Williams theorem and give each such forest a new own color.
@@ 4 Acknowledgements
The author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.

### 23 Mar 2026 v2

@@ 3 Concluding remarks
We determined the asymptotic value of $R_{h}(k;\ell)$ when $\ell=2$ and when $k\gg\ell\gg 1$. It would be interesting to extend these results, and in particular asymptotically determine the value of $R_{h}(k;\ell)$ as $\ell\geq 3$ is fixed and $k\rightarrow\infty$. It would also be nice to find an easy direct proof of an upper bound on $R_{h}(k;2)\leq ck\sqrt{\log k}$, for some, maybe large constant $c$, but avoiding the density Theorem 1.1.
To prove the lower bound on $R_{h}(k;\ell)$, one can also use the following packing result by taking $G$ being an $H$-factor, where $H$ is a graph on $k\sqrt{\log k}(1+o(1))$ vertices, density $\beta k\sqrt{\log k}$ and no $K_{k}$-minor. Pack $K_{n(1-\varepsilon)}$ with copies of $G$, give each such copy of $G$ its own color. Decompose the remaining edges of $K_{n}$ into forests using the Nash-Williams theorem and give each such forest a new own color.
@@
Note that our Theorem 1.8 nails down the value of $R_{h}(F;2)$ precisely for every graph $F$ with $\gamma$ bounded away from zero, and thus for almost all graphs on $k$ vertices with at least $k^{1+\varepsilon}$ edges for any fixed $\varepsilon>0$ [11]. However, in the sparse regime, i.e., when $F$ has only $k^{1+o(1)}$ edges, it only says that $R_{h}(F;2)=o(k\sqrt{\log k})$. Thus, it would be interesting to determine $R_{h}(F;2)$ more precisely for very sparse graphs, with, say, a constant average degree. A natural first step is to consider $R_{h}(F)$ for trees $F$ on $k$ vertices. In the case that $F=P_{k}$ is a path on $k$ vertices, we simply have $R_{h}(P_{k};2)=R(P_{k};2)=\left\lfloor\frac{3k-2}{2}\right\rfloor$ for every $k$, since any graph contains $P_{k}$ as a minor if and only if it contains $P_{k}$ as a subgraph, and since the Ramsey number of a path was determined to equal the above value by Gerencsér and Gyárfás [6]. Moreover, also for multiple colors the known results on the Ramsey numbers of paths extend to their minor-Ramsey numbers. On the other extreme, for a star $S_{k}=K_{1,k}$ on $k+1$ vertices, we show that $R_{h}(S_{k};2)$ is by a factor two smaller than $R(S_{k};2)=2k+O(1)$.
Proposition 3.2.
We have $R_{h}(S_{k};2)=k+o(k)$.
Since no graph on at most $k$ vertices contains $S_{k}$ a minor, we have $R_{h}(S_{k};2)\geq k+1$, and hence it suffices to prove the upper bound. So fix $\varepsilon>0$ arbitrarily and let us show that for every sufficiently large $k$, every red/blue-coloring of the edges of $K_{\lceil(1+\varepsilon)k\rceil}$ contains a monochromatic $S_{k}$-minor. In the following, we set $n:=\lceil(1+\varepsilon)k\rceil$. Consider any red/blue-coloring of the edges of $K_{n}$ and let $G_{r},G_{b}$ denote the spanning subgraphs of $K_{n}$ formed by the red and blue edges, respectively.
If any vertex in any of $G_{r},G_{b}$ has degree at least $k$, then our coloring contains $S_{k}$ as a monohromatic subgraph and we are done. Hence, in the following we may assume that $\Delta(G_{r}),\Delta(G_{b})<k$. Since $G_{r}$ and $G_{b}$ are complements of each other, this in particular implies that $G_{r}$ and $G_{b}$ have minimum degree at least $(n-1)-(k-1)\geq\lceil\varepsilon k\rceil=:s$ each.
Now, let $S\subseteq V(K_{n})$ be a random subset of vertices created as follows. Randomly pick vertices $v_{1},\ldots,v_{s}$ from $V(K_{n})$ with repetition, i.e. such that $v_{i}$ is uniformly random in $V(K_{n})$ independently from all other $v_{j}$. Finally, we set $S:=\{v_{1},\ldots,v_{s}\}$. We have $|S|\leq s$ and hence $|V(K_{n})\setminus S|\geq n-s=k$.
We now claim that for every sufficiently large $k$, the set $S$ simultaneously forms a dominating set in both of the graphs $G_{r},G_{b}$. Indeed, for a fixed vertex $v\in V(K_{n})$ and a fixed color $i\in\{r,b\}$, the probability that $v$ has no neighbors in $S$ in the graph $G_{i}$, i.e., that none of the vertices $v_{1},\ldots,v_{s}$ was picked from $N_{G_{i}}(v)$, is
$\left(1-\frac{d_{G_{i}}(v)}{n}\right)^{s}\leq\exp\left(-\frac{d_{G_{i}}(v)s}{n}\right)\leq\exp\left(-\frac{s^{2}}{n}\right)=\exp(-\Theta(k)).$
Hence, the probability that some vertex $v$ is not dominated by $S$ in $G_{i}$ for some color $i\in\{r,b\}$ is at most $n\cdot 2\cdot\exp(-\Theta(k))=O(k)\cdot\exp(-\Theta(k))\rightarrow 0$ for $k\rightarrow\infty$. Hence, indeed for every sufficiently large $k$, with positive probability, the set $S$ is a dominating set of both $G_{r}$ and $G_{b}$.
Now, assuming $k$ is sufficiently large, fix a set $S$ of size at most $s$ which dominates $G_{r}$ and $G_{b}$. Since $G_{r}[S]$ and $G_{b}[S]$ are complementary graphs on vertex set $S$, at least one of them, say $G_{r}[S]$, must be connected. But then we obtain $S_{k}$ as a minor of $G_{r}$ by contracting $S$ (more precisely, all edges contained in it) into a single vertex. This is the desired outcome and concludes the proof. ∎
@@ 4 Acknowledgements
The first author is grateful to Noga Alon for several very fruitful discussions and suggestions. She also thanks Jan Malte Weirich for discussions.
The second author gratefully acknowledges funding from the Ambizione grant No. 216071 of the Swiss National Science Foundation.